OpenStudy (anonymous):

Help Im Lost O.o write and simplify an expression for the perimeter of the region and write and simplify an expression for the area of the region

OpenStudy (anonymous):

Please let me know height is 2/5y+4 and parenthesis

OpenStudy (anonymous):

thats the height of the traingle i just dont know what the teacher wants me to write as an area expression and its due at 11 pm

OpenStudy (anonymous):

Please give the dimensions again and use parenthesis where needed Kris11

OpenStudy (anonymous):

their is no paranthesis it just has a picture of a triangle with the height labeled 2/5y+4 and the base=1/2y

OpenStudy (anonymous):

their is also another figure thats two rectangles

OpenStudy (anonymous):

Can you please draw the figure . It will help

OpenStudy (anonymous):

im trying but it wont let me

OpenStudy (anonymous):



OpenStudy (anonymous):

height is 2/5y+4 and base = 1/2y Area = 1/2 * base *height Area = 1/2 * 2/5y+4 * 1/2y Area = 1/(5y+4)(2y) Area = 1/ 10y^2 + 8y

OpenStudy (anonymous):

so do i write the last one?

OpenStudy (anonymous):

Yes if no parenthesis this is the answer for area

OpenStudy (anonymous):

it keep saying its wrong it only says SIMPLIFIED expression for area